OpenStudy (anonymous):

i need help i'm not good with sqrt!! Simplify the following expression sqrt -7 sqrt -105

OpenStudy (saifoo.khan):

\[\sqrt{-7 \sqrt{-105}}\]Like that?

OpenStudy (anonymous):

no they are separated like \[\sqrt{-7 \sqrt{-105}}\]

OpenStudy (saifoo.khan):

\[\sqrt{-7} \sqrt{-105}\]Like that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well, im not good at that eather.

OpenStudy (anonymous):

yeah the whole algebra thing to me is confusing

OpenStudy (anonymous):

4th grade is not at all like that, but tag is pre algebra.

OpenStudy (anonymous):

tag stands for talented and gifted.

OpenStudy (anonymous):

i am tag in math.

OpenStudy (saifoo.khan):

Just a sec.

OpenStudy (anonymous):

Oh I am in college and it's elementary and intermediate algebra I never had that in school

OpenStudy (anonymous):

ok

OpenStudy (saifoo.khan):

well, im getting... \[-7 \sqrt{15}\]

OpenStudy (saifoo.khan):

College algebra is easy as cake! :D

OpenStudy (anonymous):

not for me I never had it in high school I just recently went back to school and am taking online courses to become a nurse I really don't see how algebra relates to my career choice.

OpenStudy (saifoo.khan):

Well, it's compulsory for most of the courses offered in Online classes.

OpenStudy (anonymous):

there is no such thing as elementry algebra. but, tag homwork this week is algebra.

OpenStudy (saifoo.khan):

lol, there is that algebra as well @ash

OpenStudy (anonymous):

yeah I guess it's a requirement

OpenStudy (saifoo.khan):

Hmm, yea.. and it's good..

OpenStudy (anonymous):

if i get best grades on my tests, i will go to tag school, and then I automacatlly get to go to college.

OpenStudy (saifoo.khan):

kewl.

OpenStudy (anonymous):

but tag school is like college aredy.

OpenStudy (anonymous):

I guess it's good that I learn it since I have kids in school

OpenStudy (anonymous):

yeah.

OpenStudy (saifoo.khan):

it's different.. college algebra is math 117/116 which is taught in grades 10-11-12.

OpenStudy (anonymous):

I never took it in high school so it's all new to me

OpenStudy (anonymous):

i do college algebra beacuse I am as smart as a 10th grader.

OpenStudy (anonymous):

that is what tag means.

OpenStudy (anonymous):

thats a good thing you are that smart at math

OpenStudy (saifoo.khan):

@ash, what are u studying now?

OpenStudy (saifoo.khan):

hehe. thanks.

OpenStudy (saifoo.khan):

i have completed college algebra like 12 times! :O

OpenStudy (anonymous):

simple pre alegbra. i dont need any help though.

OpenStudy (saifoo.khan):

Cool.

OpenStudy (anonymous):

thanks. my brother wont let my brother play football until he i8s the fastest kid on his scoocer team. witch he is.

OpenStudy (anonymous):

i mean dad

OpenStudy (saifoo.khan):

Lol.

OpenStudy (saifoo.khan):

But online courses are very easy as compared to real college subjects..

OpenStudy (saifoo.khan):

Like online college math is way easier than real college math.

OpenStudy (anonymous):

ive done 8 of them and working on my 9th.

OpenStudy (saifoo.khan):

I like math.. since we have versatility to do this in out own way..

OpenStudy (anonymous):

i dont, but there is a ribbon of math in my family tree.

OpenStudy (anonymous):

my dad is a engeneer and my grandpa is a phyics teacher.

OpenStudy (saifoo.khan):

wow. My dad was a teacher, same as me. ^_~

OpenStudy (anonymous):

im thinking of being a eygptoligest.

OpenStudy (saifoo.khan):

sounds cool!!!!!o_O

OpenStudy (anonymous):

a person that studys acient eygept.

OpenStudy (saifoo.khan):

Me in future = Software Engg. :D

OpenStudy (anonymous):

oh.

OpenStudy (anonymous):

sounds exiting.

OpenStudy (saifoo.khan):

It is. :D

OpenStudy (asnaseer):

@saifoo, the answer should be:\[\sqrt{-7}\sqrt{-105}=\sqrt{(-7)*(-105)}=\sqrt{735}=\sqrt{49*15}=7\sqrt{15}\]i.e. the two minuses cancel each other out.